OpenStudy (anonymous):

find a function y=f(x) whose second derviative is y''=12x-2 at each point (x,y) on its graph and y=-x+5 is tangent to the graph at the point corresponding to x=1

zepdrix (zepdrix):

Hmm this is a neat problem :) So ummm.... y=-x+5 is tangent to the curve at x=1. The `slope` of this tangent line gives us the value of the first derivative at x=1. \(\Large\rm y'(1)=-1\) Understand how I got that information?

OpenStudy (anonymous):

yes a little thanks

zepdrix (zepdrix):

You can integrate / anti-differentiate and that data will allow you to find the unknown that shows up.

zepdrix (zepdrix):

Integrating a second time to our y function brings another constant popping up. So we need to find another piece of data we can use to figure out that constant.

zepdrix (zepdrix):

Recall that a tangent line `touches` the curve. So the tangent line and the function y will both share a coordinate pair. So plug x=1 into the tangent line, \(\Large\rm y=-(1)+5\) This will give us our coordinate pair that we can use to find the other constant term.

OpenStudy (anonymous):

so the would the final result ne f(x)= 2x^3+x^2-5x+4

zepdrix (zepdrix):

Hmm I came up with something a little different.

OpenStudy (anonymous):

what did you come up with

zepdrix (zepdrix):

\[\Large\rm y''=12x-2\]Anti-differentiating gives us:\[\Large\rm y'=6x^2-2x+c\]Plugging in our data,\[\Large\rm y'(1)=-1 \qquad\to\qquad -1=6-2+c\qquad\to\qquad c=-5\]

zepdrix (zepdrix):

So our first derivative is,\[\Large\rm y'=6x^2-2x-5\]

zepdrix (zepdrix):

Anti-differentiating again gives us,\[\Large\rm y=2x^3-x^2-5x+c_1\]

zepdrix (zepdrix):

The tangent line and our curve both pass through the point (1,4). \[\Large\rm 4=2-1-5+c_1\qquad\to\qquad c_1=8\]Something like that I think?

zepdrix (zepdrix):

\[\Large\rm y=2x^3-x^2-5x+8\]

OpenStudy (anonymous):

yeah I had did something wrong with but I get it now thank you so much :)

zepdrix (zepdrix):

cool \c:/